6-6-2 Future Interest Rates Based on the pure expectations theory, if I lock in at 4% on a two-year security, how much return do I expect on a one-year security in the third year if the other option is to lock in a yield of 3.5% for the next three years? 6-6-3 Future Interest Rates -2nd Example Suppose that you can lock into a five-year security with a yield of 4%. If the threeyear rate is 3.5%, what is the expected two-year rate in three years (for years four and five) based on the pure expectations theory? 6-7 Macro Factors The Federal Reserve decides to greatly increase the money supply. What is the expected effect on interest rates? Lower interest rates
Answer #1

on a two year security I earn = 4% per annum for 2years =r02

on a three year security I earn = 3.5% per annum for3 years = r03

on a one year security I earn = x% per annum in 3rdyears = E23

E23 = (1+r03)3 / (1+r02)2 - 1

E23 = (1+0.035)3/ (1+.04)2 - 1

E23 = 2.507~ 2.51%

as per Per pure expectation theory I expect on a one-year security in the third year is 2.51%

explanation:-

In two year I earn = 1.04*1.04 = 1.0816 or 8.16%

in three year I earn = 1.035*1.035*1.035 =1.1087 or 10.87%

i.e. upto 3 year if have to earn 10.87% ,then in 3rd my interest rate should be such that 2 year interest and 3rd year intt combines to 10.87%

1.0816*1.02507 = 1.1087 or 10.87%

according to the HOMEWORKLIB RULES i can solve only 1 please post another part seperately

thank you let me know if you have any doubt